Describe how you would prepare a buffer solution from NaH2PO4 and Na2HPO4 to have a pH of 7.5.
1
Expert's answer
2020-11-12T06:36:19-0500

pH = pKa + log ((Na2HPO4)/(NaH2PO4))

7.50 = 7.21 + log ((Na2HPO4)/(NaH2PO4))

log ((Na2HPO4)/(NaH2PO4)) = 7.50 - 7.21 = 0.29

(Na2HPO4)/(NaH2PO4) = 1.3364

this is the mole ratio

let's say we take

1 mole of NaH2PO4 (1 x 119.98 = 119.98 g)

and 1.3364 mole of Na2HPO4 (1.3364 x 141.96 = 189.72 g)

dissolve the salts in enough water to make 1 L solution